OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{prove\show that the equation has no solution over integers} \hspace{.33em}\\~\\ (x^3 + 7 y^2 + 5)=0 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\(\large\begin{align} \color{black}{ \normalsize \text{one method is the } \pmod{7} \hspace{.33em}\\~\\ \normalsize \text{i would like to know if their are other methods as well .}\hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

we can prove a much stronger statement : \[\large x^3 + 7y^{n} + 5 = 0\] has no solutions over intgers \(x, y\) and whole number \(n\)

OpenStudy (mathmath333):

yes beacause that goes way with 7 ?

OpenStudy (mathmath333):

\(y^n\)

ganeshie8 (ganeshie8):

Yes.. below is the most simplified form that the question can be framed : Prove that \(a^3+5\) is not divisible by 7 for all integers \(a\)

ganeshie8 (ganeshie8):

show me your solution xD

OpenStudy (mathmath333):

\((1)^3 \equiv1 \pmod{7} \\ (2)^3 \equiv1 \pmod{7}\\ (3)^3 \equiv6 \pmod{7}\\ (4)^3 \equiv1 \pmod{7}\\ 5^3 \equiv6 \pmod{7}\\ 6^3 \equiv 6 \pmod{7}\\ \)

ganeshie8 (ganeshie8):

that works because 7 is small! how do we prove below : \(\large a^3 + 5 \) is not diviisble by 43 for all integers \(a\) :P

OpenStudy (mathmath333):

oh let me think

ganeshie8 (ganeshie8):

mod 43 works, the solution will be lengthy thats all

OpenStudy (mathmath333):

how did u worked out \(43 \)?

ganeshie8 (ganeshie8):

You want to show \(a^3+5\) is not divisible by \(43\) for all integers \(a\) simply consider 43 cases for \(a\) : {43k, 43k+1, ..., 43k+42}

OpenStudy (mathmath333):

ohk i get that

OpenStudy (mathmath333):

but that too lengthy

ganeshie8 (ganeshie8):

By binomial theorem \((43k+l)^3 = 43X + l^3 \equiv l^3 \pmod{43}\)

ganeshie8 (ganeshie8):

Yeah, that method is called division algorithm you can prove almost ALL divisibility problems using that method, but as you can see the proof is very boring..

OpenStudy (mathmath333):

does the primitive root can be used to quickly work out \(mod ~43\)

OpenStudy (mathmath333):

if not that seems to be programming question

ganeshie8 (ganeshie8):

it can be done in one line using primitive roots if you have a table for indices of any primitive root of 43

ganeshie8 (ganeshie8):

start by finding out a primitive root of 43

OpenStudy (mathmath333):

3,5,12,18 are one of those

ganeshie8 (ganeshie8):

yes, lets use 3 as base of our logarithm as it is the smallest number.. so it will be easy to work out powers..

OpenStudy (mathmath333):

if we use 3 soo we not need to use other primitive roots of 43 ?

ganeshie8 (ganeshie8):

one primitive root is enough..

OpenStudy (mathmath333):

oh that cool!

ganeshie8 (ganeshie8):

we want to show below congruence has no solutions : \[a^3+5 \equiv 0 \pmod{43}\]

OpenStudy (mathmath333):

we take log both sides

ganeshie8 (ganeshie8):

which is same as \[a^3 \equiv -5 \pmod{43}\]

ganeshie8 (ganeshie8):

which is same as \[a^3 \equiv 38 \pmod{43}\] yes now take log base 3 both sides

OpenStudy (mathmath333):

\(\large\begin{align} \color{black}{ 3ind~a \equiv ind 38 \pmod{43}\hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

since this is a special logarithm, we use a different notation : \(\text{ind}_3\)

ganeshie8 (ganeshie8):

looks good!

ganeshie8 (ganeshie8):

lets pause a bit for few minutes

OpenStudy (mathmath333):

ok

OpenStudy (mathmath333):

yes wait

ganeshie8 (ganeshie8):

sorry not that, one sec..

OpenStudy (mathmath333):

u said thar it should be \(p-1\)so \(43-1=42\)

ganeshie8 (ganeshie8):

Exactly! we should end up with below congruence to solve : \[\large 3\text{ind}_3\ a \equiv \text{ind}_3 38 \pmod{\color{Red}{42}}\]

ganeshie8 (ganeshie8):

our goal is to prove that above congruence has no solutions thats same as proving that the congrunce \(a^3 = 38 \pmod {43}\) has no solutions because we have just taken log both sides..

ganeshie8 (ganeshie8):

lets pause a bit now and see how to solve linear congruences in general

ganeshie8 (ganeshie8):

do you know how to solve below congrunce : \[3x \equiv 22 \pmod{42}\] ?

OpenStudy (mathmath333):

wait let me see

ganeshie8 (ganeshie8):

trial and error is a pain there is really a neat way to work this..

OpenStudy (mathmath333):

do i have to put all no upto \(41\) for \(x\)

ganeshie8 (ganeshie8):

yes that works, but lets see if we can work it without doing all that work..

ganeshie8 (ganeshie8):

\[3x \equiv 22 \pmod{42}\] is same as \[3x = 42k + 22\] for all integers \(k\), yes ?

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

\[3x = 42k+22\] \[3x -42k=22\] \[3(x -14k)=22\] why does this equation makes no sense ?

OpenStudy (mathmath333):

may be it has no solution

ganeshie8 (ganeshie8):

yes, why ?

OpenStudy (mathmath333):

the 3 should cancel out i think

ganeshie8 (ganeshie8):

\[3(x -14k)=22\] this equation is saying \(3\) times some integer equals \(22\) which is not possible. so the congrunce \(3x\equiv 22 \pmod{42}\) has NO SOLUTIONS.

OpenStudy (mathmath333):

so is hypithesis of cancelling out 3 for integers right ?

OpenStudy (mathmath333):

*hypothesis

ganeshie8 (ganeshie8):

3*t = 22 can you find an integer "t" that satisfies above equaiton ?

OpenStudy (mathmath333):

lol no3*7=21,3*8=24

ganeshie8 (ganeshie8):

\[3(x -14k)=22\] has no solutions in integers for the same reason

OpenStudy (mathmath333):

ok

ganeshie8 (ganeshie8):

because 3 is not a factor of 22

ganeshie8 (ganeshie8):

as simple as that ^

OpenStudy (mathmath333):

so the indices not solvable

ganeshie8 (ganeshie8):

Yes!

OpenStudy (mathmath333):

oh

ganeshie8 (ganeshie8):

\[\large 3\text{ind}_3\ a \equiv \text{ind}_3 38 \pmod{\color{Red}{42}}\]

ganeshie8 (ganeshie8):

if you prepare the indices table, you will see that \(\text{ind}_3 38 = 22\)

OpenStudy (mathmath333):

so tha'ts why it has no solution ?

ganeshie8 (ganeshie8):

so the congruence becomes \[\large 3\text{ind}_3\ a \equiv 22\pmod{\color{Red}{42}}\]

OpenStudy (mathmath333):

ok

ganeshie8 (ganeshie8):

we have just worked that this congruence has NO solutions

ganeshie8 (ganeshie8):

it will take some time to get use to these logarithms and congruences.. i see you're very good with congruence equaitons :)

OpenStudy (mathmath333):

for integers

ganeshie8 (ganeshie8):

\[\large 3\text{ind}_3\ a \equiv 22\pmod{\color{Red}{42}}\] has no solution shows that \(a^3 + 5\) is not divisible by 43

OpenStudy (mathmath333):

sry this one \(a^2+5 \pmod {43})\)

ganeshie8 (ganeshie8):

yes

OpenStudy (mathmath333):

ohk thats not long :)

OpenStudy (mathmath333):

than division algo

ganeshie8 (ganeshie8):

yes its not long because we have used the indices table : \(\text{ind}_3 38 = 22\) working it manually might take some time :P

ganeshie8 (ganeshie8):

let me put the complete solution in one reply

ganeshie8 (ganeshie8):

Prove \(a^3+5\) is not divisible by \(43\) for all integers \(a\) Proof : \[a^3 \equiv -5 \pmod{43}\] \[a^3 \equiv 38 \pmod{43}\] take log base 3 both sides \[3\text{ind}_3a \equiv \text{ind}_3 38 \pmod{42}\] but we know from indices table that \(\text{ind}_3 38 = 4\) : http://www.wolframalpha.com/input/?i=3%5E4+mod+43 so the congrunce becomes \[3\text{ind}_3a \equiv 4 \pmod{42}\] which is same as \[3\text{ind}_3a = 42k+ 4\] which is same as \[3(\text{ind}_3a -14k) = 4\] which is impossible because 3 is not a factor of 4.

OpenStudy (mathmath333):

\(ind_3 38 \pmod {42}\) =\(3^x \mod 43\)

ganeshie8 (ganeshie8):

indices are exactly same as logarithms

ganeshie8 (ganeshie8):

\(\log_3 81 = 4\) because \(81 = 3^4\) \(\text{ind}_3 38 = 4\) because \(38\equiv 3^4 \pmod{43}\)

OpenStudy (mathmath333):

so if i need to find \(ind_3 38 \pmod {42}\) then i need to find this ? \(3^x \pmod {43}=38\)

ganeshie8 (ganeshie8):

tell me how would you find below : \[\log_2 1024\] ?

OpenStudy (mathmath333):

\(1024=2^{10}\)

OpenStudy (mathmath333):

so its \(10\) ?

ganeshie8 (ganeshie8):

yes, but how do you know 1024 = 2^10 ?

ganeshie8 (ganeshie8):

may be you worked out all the powers up to 10 ?

OpenStudy (mathmath333):

that is intuition !

ganeshie8 (ganeshie8):

2^1 = 2 2^2 = 4 2^3 = 8 ... 2^10 = 1024 so \(\large \log_2 1024 = 10\)

ganeshie8 (ganeshie8):

similarly lets work out \(\large \text{ind}_3 38\) in modulo \(43\)

ganeshie8 (ganeshie8):

you want to solve below : \[38 \equiv 3^x \pmod{43}\]

ganeshie8 (ganeshie8):

x = 4 works because 3^4 = 81 and 38 - 81 = -43 which is divisible by 43

OpenStudy (mathmath333):

ok that seems cool

ganeshie8 (ganeshie8):

that means \(38 \equiv 3^4 \pmod{43}\) so \(\text{ind}_3 38 = 4\)

OpenStudy (mathmath333):

this whole stuff is cool

ganeshie8 (ganeshie8):

i just want you know that this is an undergraduate course and students sepnd 1 full semister on these... so you should be proud of yourself for making sense of these in just couple of hours... :)

OpenStudy (mathmath333):

O: :O